OpenStudy (anonymous):

if tan =-1/3 and 90 degrees is less than equal to theta less than equal to 180 degrees, what is the exact Value of sec theta. I KNOW THE ANSWER I JUST NEED HELP

OpenStudy (noelgreco):

Which quadrant?

OpenStudy (anonymous):

not really sure on this one

OpenStudy (anonymous):

not to good with tan

OpenStudy (noelgreco):

If\[90 \le \theta \le 180\] which quadrant are you in? Don't worry about the trig function yet.

OpenStudy (anonymous):

2nd quadrant

OpenStudy (noelgreco):

Good. Now draw a terminal side in the 2nd quadrant like before.

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):



OpenStudy (noelgreco):

 Remember tan x = opp/adj